OpenStudy (anonymous):

e^(2x + 4) - 7 = 0 solving for x log( e^(2x + 4))=log(7) 2x+4=log(7) x=(log(7)-4) /2

OpenStudy (anonymous):

is this right or not?

OpenStudy (anonymous):

it is right

OpenStudy (anonymous):

cause i tried this answer in my quiz and its wrong

OpenStudy (anonymous):

but it also says in the question "Solve each equation for x. Write the exact value: do not write an approximation."

OpenStudy (anonymous):

does that mean just a value?

ganeshie8 (ganeshie8):

try \(\large \frac{1}{2}(\ln 7 - 4)\)

ganeshie8 (ganeshie8):

the online grader is much dumber than you think it is.. you need to try all possible variations sometimes...

OpenStudy (anonymous):

i only get 3 tries and i used up 1

ganeshie8 (ganeshie8):

then try try this : \(\large \frac{1}{2}(\ln( 7) - 4)\)

OpenStudy (anonymous):

it also says "Note: To see what your typed text represents, click on the picture of an eye to the right of the answer box. However, if you use the allowed exp(x) for e^x, the eye will still display it as exp(x) rather than as e^x.

OpenStudy (anonymous):

does that mean i have to represent the answer in exponential form?

ganeshie8 (ganeshie8):

i don't think its that stupid.. pray ur best gods and try above form >.<

ganeshie8 (ganeshie8):

but wait a sec, why are you doing this algebra problem in calculus course ?

OpenStudy (anonymous):

lol i'll just ask my lecturer 2moro, can't afford to get it wrong

OpenStudy (anonymous):

i think they r just trying to get us to learn the log rules before we do integration with logs

OpenStudy (anonymous):

also how would i do the above question?

ganeshie8 (ganeshie8):

differentiate implicitly both sides

OpenStudy (anonymous):

no the first one isn't asking for derivative

ganeshie8 (ganeshie8):

\[\large e^{x^7y} = x+y\]

ganeshie8 (ganeshie8):

differentiate both sides with respect to x, implicitly

ganeshie8 (ganeshie8):

\[\large \left( e^{x^7y} \right)'= \left(x+y\right)'\]

ganeshie8 (ganeshie8):

\[\large e^{x^7y} \left( x^7y \right)'= 1 + y'\]

ganeshie8 (ganeshie8):

can you take it home from here ?

OpenStudy (anonymous):

i have forgot how to do implicit differentiation ,let me look back at my notes

ganeshie8 (ganeshie8):

you just differentiate as you would differentiate product of two different functions of x : f(x)*g(x)

ganeshie8 (ganeshie8):

y = g(x) here, so u need to apply chain rule

ganeshie8 (ganeshie8):

is the first step clear ?

ganeshie8 (ganeshie8):

derivative of e^x is just e^x so derivative of \(\large e^{x^7y} \) is just \(\large e^{x^7y} \) not exactly, you need to use chain rule because the exponent is not just x

ganeshie8 (ganeshie8):

\(\large \left(e^{x^7y} \right)' = e^{x^7y} \left(x^7y \right)' \)

ganeshie8 (ganeshie8):

if that makes any sense..

OpenStudy (anonymous):

but how do u (x^7y)'

ganeshie8 (ganeshie8):

we use product rule

ganeshie8 (ganeshie8):

\[\large \left(fg\right)' = f'g + fg'\]

ganeshie8 (ganeshie8):

\[\large \left(x^7y\right)' = \left(x^7\right)' y + x^7 \left(y\right)'\]

OpenStudy (anonymous):

7x^6y _x^7(y)'

ganeshie8 (ganeshie8):

from where did _ came from hmm

ganeshie8 (ganeshie8):

\[\large \left(x^7y\right)' = \left(x^7\right)' y + x^7 \left(y\right)' = 7x^6y + x^7y'\] right ?

ganeshie8 (ganeshie8):

\[\large e^{x^7y} \left( x^7y \right)'= 1 + y'\] \[\large e^{x^7y} \left(7x^6y+x^7y'\right)= 1 + y'\]

ganeshie8 (ganeshie8):

isolate \(\large y'\)

OpenStudy (anonymous):

how did u get 1+y'

ganeshie8 (ganeshie8):

thats the derivative of right hand side : x + y

OpenStudy (anonymous):

sorry i was trying to find my notes on this

OpenStudy (anonymous):

i tried to isolate y', but failed how would u of it?

OpenStudy (anonymous):

@Kainui

ganeshie8 (ganeshie8):

\(\large e^{x^7y} \left(7x^6y+x^7y'\right)= 1 + y' \) \(\large e^{x^7y} \cdot 7x^6y+ e^{x^7y} \cdot x^7y' = 1 + y' \) \(\large e^{x^7y} \cdot 7x^6y - 1 = y'\left(1 - e^{x^7y} \cdot x^7 \right) \) \(\large \dfrac{e^{x^7y} \cdot 7x^6y - 1 }{1 - e^{x^7y}\cdot x^7} = y'\)

ganeshie8 (ganeshie8):

something like that maybe ? if grader cries, you may substitute \(\large e^{x^7y}\) by \(\large x+y\) so that : \(\large \dfrac{(x+y) \cdot 7x^6y - 1 }{1 - (x+y)\cdot x^7} = y'\)

OpenStudy (anonymous):

why is e^x^7y substituted with (x+y)?